# Towards automatizing detection and quantification of intestinal metaplasia: a multi-expert comparative study

### v1

@@ Discussion
This paper has investigated the role of state-of-the-art deep learning models to detect and quantify intestinal metaplasia. Evaluation of these models has demonstrated remarkable performance for computing stroma, glands and the percentage of actual intestinal metaplasia. However, estimation of the OLGIM score is still a bottleneck since pathologists show high variability at estimating the intestinal metaplasia percentage, a result consistent with publications which have shown and discussed OLGIM limitations [16, 18, 20]. As inferred from the presented results, decision uncertainty is still too high as to conclude that OLGIM should be the only element at managing and assigning a level of risk.
Several studies have investigated OLGIM limitations, among them the most important is the difficulty to report, i.e., many pathologists worldwide hardly report OLGIM stages since they are aware of the conflict introduced by the typical disagreement [16], a moderate agreement is widely acknowledged [56, 54]. An inevitable criticism to OLGIM score is that this is by nature biased since intermediate stages are obtained with a higher number of combinations and therefore these stages are the more frequent, in other words dices are charged and as a result the confidence in this assessment ends up by being undermined.
Recent studies have demonstrated the effectiveness of deep learning models at classifying and grading pathologies in H&E images [24, 27, 32]. Typically, the pathologist knowledge, the ground truth [57], consists in a set of delineated regions which serve to train models that closely replicate these visual annotations. However, the variable biology along with a subjective judgment, inevitably introduces additional biases. As a general rule, artificial intelligence models are good if data is available in the needed quantity and quality [58]. However, annotations are an actual bottleneck since they are typically closed regions which include not only metaplasia, but in a great proportion stroma and normal glands [59]. The agreements obtained here between experts and model predictions suggest further limitations: while models perform well at identifying regions with intestinal metaplasia and at estimating percentage, agreement with experts remains low, as reported in other publications [54]. Nevertheless, the larger expert variability at estimating percentages when comparing with models is likely compensated by experience, a probable explanation of the OLGIM overestimation herein observed, and a result again consistent with studies in the literature about OLGIM limitations [16, 18].
@@
Despite the multifocal nature of the intestinal metaplasia, i.e. spreading along stomach regions, estimation of this condition considers this phenomenon as a binary problem: presence or absence. Both intestinal metaplasia extension and topographic distribution have been at the very base of risk estimation to progress, yet Correa et al. had already questioned this description because intestinal metaplasia is a mixture of morphological expressions of the different mucin enzymes associated to two main subtypes, complete and incomplete [10]. While complete metaplasia is characterized by the tissue resemblance to the small intestine with enterocytes displaying the brush border, the incomplete metaplasia is close to the colonic epithelium with multiple and irregular intra-cytoplasmic mucin droplets of variable size. Current estimation of metaplasia is still far from quantifying this level of detail, but clearly these AI tools may build a continuous description of the phenomenon instead of the limited binary characterization so far applied. In the present study deep learning models ended up by describing metaplasia as a binary phenomenon, aiming to drive any pattern to any of these two classes but with a variable uncertainty introduced by the presence of different morphological expressions of metaplasia. Consequently, these models could quantify not only the metaplasia presence/absence but also a continuous characterization of the lesion profiles, a necessary step towards a reproducible quantification and a personalized risk estimation.
@@ Conclusion
Machine learning models demonstrate to accurately quantify the intestinal metaplasia proportion. Likewise, the model is clearly reproducible and introduces an objective factor in the chain of events leading to estimate the risk.
@@
Conceptualization: FC, ACR, FG and ER; Data curation: FC, MC, AS, JV and JQ; Methodology: FC, MC, ACR, FG and ER; Formal analysis: FC, ACR, FG and ER; Writing - original draft preparation: FC and ER; Writing - review and editing: FC, ACR, FG and ER; Funding acquisition: ACR and ER; Resources: ACR, FG and ER; Supervision: FG and ER; Project administration: ER. All authors have read and agreed to the published version of the manuscript.
@@ Acknowledgements
Internal and external data are publicly available at Harvard Dataverse [60].

### v3

@@ Discussion
This paper has investigated the role of state-of-the-art deep learning models to detect and quantify intestinal metaplasia (IM) on digitized biopsy samples from patients and volunteers. Evaluation of these models demonstrated relatively accurate performance for computing stroma, glands, and the percentage of IM, with F1 = $0.80$ and AUC = $0.91$. Comparison of OLGIM scoring, pathologist assessments, and deep learning predictions highlighted the inherent subjectivity of visual estimation, with only moderate inter-observer agreement (Fleiss’ Kappa = $0.31$ for the antrum and $0.41$ for the corpus), consistent with previously reported values in the literature [54] (Fleiss’ Kappa = $0.48$). These findings confirm both the limitations of current risk systems, limiting the certainty of the medical decision and patient confidence, and the potential of AI tools to complement pathologist-based evaluation by serving as reproducible “second readers”.
Several studies have investigated OLGIM limitations, and many pathologists worldwide hardly report OLGIM stages since they are aware of the conflict introduced by the typical disagreement [16], indeed, a moderate agreement is widely acknowledged [55, 54]. An inevitable criticism of the OLGIM score is that it is inherently biased, that is, intermediate stages are obtained with a higher number of combinations and therefore these stages are more frequent, in other words “dices are charged” and confidence therefore is undermined. Additionally, pathologists tended to overestimate the extent of the lesion [18], typically assigning higher percentages than the model. In the present investigation, the analysis was restricted to cross-sectional biopsies, limiting thereby the impact of the present study. Moreover, annotations represented a bottleneck, since delineated regions often included not only metaplasia but also stroma and normal glands, introducing bias [56]. The multifocal nature of IM and its patch-like distribution further complicates assessment, as even standardized sampling with the five-biopsy protocol may fail to capture the full extent of lesions, affecting both manual and automated quantification. This work did not distinguish between complete and incomplete IM subtypes, which are morphologically and prognostically distinct, due to the lack of annotations at this level of detail. Finally, the participating pathologists were recently graduated, which may partly explain the relatively low inter-observer agreement observed.
It is important to recall that, in this work, OLGIM staging relied exclusively on IM since atrophy was difficult to evaluate, and even among experienced pathologists it has shown poor inter-observer reproducibility [19]. In contrast, IM has been recognized as the most reliable and prognostic component of OLGIM staging [57]. This methodological choice therefore prioritizes objectivity and reproducibility. Recent studies have further demonstrated the effectiveness of deep learning models at classifying and grading pathologies in H&E images [24, 27, 32]. Typically, pathologist knowledge, “the ground truth”[58], consists of delineated regions that train models to replicate visual annotations. However, variable biology along with subjective judgment inevitably introduce additional biases.
@@
Despite the multifocal nature of IM, i.e. spreading along stomach regions, estimation of this condition is often reduced to a binary problem: presence or absence. Both IM extension and topographic distribution have been at the very base of risk estimation for progression. Correa et al. highlighted that IM comprises two main subtypes, complete and incomplete, associated with distinct morphological expressions of mucin enzymes. Complete IM resembles the small intestine, with enterocytes displaying a brush border, whereas incomplete IM is closer to colonic epithelium with irregular intra-cytoplasmic mucin droplets of variable size. In the present study, we did not differentiate between these subtypes because annotations at this level of detail were not available, and our analysis focused on overall IM quantification in line with OLGIM staging. We acknowledge this as a limitation, and future research with dedicated annotations and tailored AI methods may enable reliable distinction of IM subtypes, moving beyond binary characterization and towards a more continuous description of the phenomenon, which is a necessary step for reproducible quantification and personalized risk estimation.
In summary, deep learning-based quantification of IM demonstrates robust performance, reproducibility, and potential clinical relevance. While expert variability, annotation bias, and sampling constraints remain significant challenges, computational pathology offers a pathway for systematic and objective evaluation. With further validation in multi-center cohorts, follow-up studies, and the integration of IM subtyping, AI-based tools could transform gastric cancer prevention strategies, moving from subjective visual estimation to reproducible, data-driven, and patient-tailored risk assessment.
@@ Conclusion
Deep learning models can reliably detect and quantify intestinal metaplasia with high performance, offering consistent estimates that may reduce the subjectivity of expert-dependent visual assessment. Further validation in larger, multi-center cohorts and the inclusion of complete and incomplete subtypes will be essential to capture the full morphological spectrum. In the long term, reproducible AI-based quantification of precancerous lesions has the potential to enhance gastric cancer risk stratification and guide more objective, personalized clinical decisions.
@@
Conceptualization: FC, ACR, FAG, SEV and ER; Data curation: FC, MC, AS, JV and JQ; Methodology: FC, MC, ACR, FAG and ER; Formal analysis: FC, ACR, FAG, SEV and ER; Writing - original draft preparation: FC and ER; Writing - review and editing: FC, ACR, FAG, SEV and ER; Funding acquisition: ACR and ER; Resources: ABU, MCB, YYC, ACR, FAG and ER; Supervision: FAG, SEV and ER; Project administration: ER. All authors have read and agreed to the published version of the manuscript.
@@ Acknowledgements
The collection of internal gastric samples used in this study were obtained from the primary research project entitled Investigación de la prevalencia de lesiones precursoras de malignidad gástrica y efecto de la erradicación de Helicobacter pylori colon prevención primaria de cáncer gástrico en el Departamento de Nariño. This project was approved under Agreement No. 057 of 2017 by the Collegiate Body for Administration and Decision (OCAD Pacífico), and was funded by the Science, Technology, and Innovation Fund of the General System of Royalties/Government of Nariño. Internal and external data are publicly available at Harvard Dataverse [59].